# Quantum Computing in Wireless Communications and Networking: A Tutorial-cum-Survey

### v1

@@ II-F Quantum-inspired Algorithms
The origin of quantum-inspired algorithms can be attributed to the integration of QC concepts into conventional algorithms. Typically, algorithms in this class are described as running on classical computers, with some steps imitating properties unique to the quantum realm.
A subclass of these algorithms mentions the implementation on a quantum computer. Consider a QRL algorithm, which denotes the state transition via the measurements of an observable [86]. The previous state will collapse to one of the observable eigenstates (next state) associated with an action chosen probability by the agent. In theory, this seems plausible. Nevertheless, for a complex application scenario, physically representing and implementing an observable at each step could be an issue. Moreover, the no-cloning principle makes it unclear how to obtain sufficient independent quantum state copies for the statics of measurement results. Therefore, the algorithm is classified as quantum-inspired RL [101].
The second subclass of these algorithms explicitly states that they are running classically. A well-known example is Tang’s work [102], in which a binary tree-like data structure is proposed to satisfy quantum state preparation assumptions. A classical algorithm exploiting this structure can output an $\mathcal{l}^{2}$-norm sample from a rank-k approximation of a matrix in nearly polynomial time. The third category is named quantum-inspired metaheuristic algorithm. As the name suggests, these algorithms modify steps of classical metaheuristic algorithms inspired by quantum principles so as to improve performance [103]. There are many related research works, including quantum ant colony algorithm [104], quantum evolutionary algorithm [105], QGA [106], QPSO [107], to solve well-known fundamental NP-hard problems such as knapsack, traveling salesman to practical issues involving clustering, project scheduling, etc.
@@
The modifications to quantum-inspired algorithms based on QC principles follow a similar pattern: utilizing qubits, or n-dimensional normalized vectors, to map and represent the original data, such as chromosomes in genetic algorithms; employing rotation gates or matrices to evolve qubits for exploration; using a random number generator to generate new feasible solutions, simulating quantum measurements collapsing to ground states; integrating entangled states in qubit representation, so that when creating a new population by measuring the qubit string (solution), the entangled qubit pairs in the solution diversify the data and help in avoiding stagnation of local optima.
@@
Quantum-inspired metaheuristic algorithms enhance optimization performance by leveraging QC concepts to accelerate convergence, enhance exploration and exploitation, and find the best solutions more effectively. However, most published work still requires in-depth complexity analysis to justify performance benefits. On the other hand, how to appropriately embed QC into optimization workflow, be compatible with the metaheuristic algorithms, and not reduce the algorithm’s speed is still a concern.
@@ II-G Common QC Platforms and Libraries
• Superconducting: Superconducting is the most popular technical route to produce QC hardware. This technique uses superconducting circuits to generate and manipulate qubits and employs microwave pulses to control quantum gates. IBM Quantum and Google Sycamore quantum computers are both grounded in this technique route.
-Pros: fast response, scalable, and seamlessly integrated with existing microelectronics technology. Moreover, these circuits can be mass-produced and integrated at the chip level.
-Cons: Superconducting qubits are susceptible to environmental perturbations, including thermal noise, electromagnetic noise, etc., resulting in shorter coherent time.
-Challenges: maintain high fidelity while scaling.
• Ion Trap: Ion trap QC is a relatively mature technique that traps ions with light and magnetic fields and manipulates them with lasers. Typical products include IonQ Aria and Honeywell System Model H.
-Pros: High stability, long coherent time, and high quantum gate fidelity, making them well-suited for QC. Moreover, the ion trap technique can practically manipulate all qubits, unlike the superconducting technique, which requires a long chain of operations on non-adjacent qubits.
-Cons: Cumbersome physical manipulation for ion traps, high optical stability requirements, and complex laser operating systems.
-Challenges: complex to scale.
• Optical QC: Using photons as information carriers. Xanadu Borealis is one representative machine.
-Pros: less affected by environmental noise and have relatively mature techniques for generating, manipulating, and detecting single photons.
-Cons: require precise optical operations, reliable single-photon sources, and efficient single-photon detectors.
Challenges: implement quantum logic gates with high efficiency and high fidelity, as well as large-scale optical QC.
• Topological QC: Topological QC relies on emergent quasiparticles, called Majorana zero-modes, that are robust against perturbations. Microsoft has been engaged in this technique route.
-Pros: more stable qubits, resistant to many types of error and noise.
-Cons: manufacturing and controlling Majorana zero-modes is complex and technically challenging.
Challenges: have not reliably fabricated and detected Majorana zero-modes in experiments.
@@
• Qiskit [108]: Qiskit is an open-source software development kit (SDK) developed by IBM. It provides tools for programming on IBM quantum computers or simulators on a local computer. Qiskit uses the Python programming language. Its official documentation provides a wealth of tutorials and case studies for QC practitioners to learn, practice, and conduct research. It is worth mentioning that Qiskit, combined with IBM’s quantum computers, provides an excellent QC ecosystem for researchers worldwide. As a software project, Qiskit has developed rapidly and constantly integrated new achievements. For example, Qiskit Machine Learning has been incorporated into the Qiskit platform, providing various tools and algorithms for building and training quantum machine learning models, which can handle numerous tasks such as classification, regression, and RL.
@@
• Cirq & TensorFlow Quantum [109]: Cirq is developed by the Google AI Quantum Team. It provides a powerful simulator and also supports programming for specific quantum processors. It has two distinctive components: OpenFermion, mainly applied to quantum chemistry, and TensorFlow Quantum (TFQ), an open-source machine learning library based on TensorFlow. It introduces two primitive data types: Quantum circuit batch, which represents the quantum circuit with varied sizes, and Pauli sum batch, which represents a linear combination of tensor products of Pauli operators. Exploiting these two data types, TFQ can efficiently sample the circuit batch’s output distribution, compute the expected values of Pauli sum and circuit batches, calculate the gradient, and simulate quantum circuits and quantum state batches. TFQ practitioners can not only leverage TensorFlow’s functionality, such as automatic differentiation and machine learning tools, but can also utilize quantum circuit simulators and connect quantum hardware. Yet, since TFQ is closely related to TensorFlow and Cirq, it faces a compatible challenge with other QC libraries or frameworks.
@@
• PennyLane [110]: PennyLane is a Python-based quantum machine learning library that seamlessly integrates classical machine learning algorithms with QC using automatic differentiation. It operates independently of any specific underlying framework and smoothly interfaces with both classical machine learning frameworks (like PyTorch and TensorFlow) and QC frameworks (such as Qiskit and Cirq). Utilizing a plugin system, PennyLane enables users to connect to various quantum hardware and simulators without the need for custom code for each platform. Its libraries and modules encompass Quantum-specific operations for evaluating and optimizing quantum circuits, Quantum nodes for encapsulating quantum circuits usable in machine learning, plugins facilitating the execution of quantum circuits on real hardware or simulators, and Quantum Machine Learning components for integrating quantum circuits with classical machine learning models, including neural networks, to develop hybrid quantum-classical machine learning models. PennyLane is highly cross-platform compatible, has automatic differentiation quantum nodes for training machine learning models, and provides user-friendly interfaces for quantum algorithms and applications that are easy to understand and implement. On the other hand, its performance and functionality strongly depend on the stability and quality of the plugin system. Besides, the PennyLane community is relatively small and needs more supporting resources than other tools grounded in traditional machine learning frameworks.
@@
• Torch Quantum (QuantumNAS) [111]: Torch Quantum, developed by a consortium including MIT, Yale, and the University of Texas at Austin, integrates seamlessly with PyTorch, adding several key components: a Quantum State Vector Simulator optimized for GPU acceleration, Quantum Circuit Modules enabling the definition and manipulation of PQCs akin to classical neural networks, Quantum Optimization facilitating parameter optimization using classical optimization algorithms like gradient descent, and an Interface to real Quantum devices. This extension of PyTorch allows developers to construct quantum programs using familiar tools and interfaces while benefiting from automatic differentiation for training QNNs. Additionally, the High-Performance Computing State Vector Simulator supports large-scale simulations on GPUs. However, it may not match the performance of specialized QC platforms like IBM Quantum and Rigetti due to its reliance on PyTorch optimization.
@@ III-A Overview
Figure 12: Different network scenarios in wireless communications.
@@ III-B1 Overview
Power allocation within wireless communication [112] involves the strategic distribution of available power resources among the diverse components of a communication system, such as base stations and user devices. The primary objective is to optimize power utilization to achieve communication that is both reliable and efficient, taking into account factors such as signal quality, network coverage, interference management, and energy consumption [113], which is also an important issue in 5G/6G cellular networks [114]. In more straightforward terms, power allocation revolves around determining the appropriate amount of power assigned to each transmitter or device in a wireless network to ensure effective communication.
Power allocation is usually implemented at the physical layer, involving only nodes within one same communication cell, optimizing the transmission power of signals to maximize transmission efficiency or minimize bit error rate. For example, power optimization allocation schemes with relay selection have been studied [115]. By selecting appropriate relay nodes and allocating corresponding power, the energy efficiency is maximized. Power allocation in wireless communication systems is primarily aimed at mitigating issues arising from the considerable variation in received signal strength due to varying distances between transmitters and receivers. This variance can lead to degradation in system performance, necessitating power control to balance the transmission power.
TABLE III: Summary of QC Works on Power Allocations in Wireless Networks
@@
[117] Primary users and secondary users competing for spectral resources Employ Bernstein approximation to address probability constraints and co-evolutionary theory to handle large-scale problems. Maximize network throughput
[115] One-hop cooperative relay between BS and multiple users in a NOMA system Apply QPSO to maximize QoS from the perspective of MIMO Maximize computation efficiency and energy efficiency
Overlay networks [118] Primary users and one primary BS; Cognitive users and one cognitive BS Formulate resource scheduling in a two-layer wireless network, and propose a model to describe outage probability Minimize system outage probability
[119] K users and N sub-carries in OFDMA system: two heterogeneous services, minimum-rate guarantee and best effort services Divide the problem: channel assignment given power allocation, power allocation given channel assignment; transform it into a concave optimization and solved by Dinkelbach method Maximize energy efficiency
@@
TABLE IV: Details of QC Algorithms for Power Allocation
@@
In this section, we undertake an investigation of existing QC-driven approaches along two fundamental dimensions: the complexity of network scenarios and the categories of QC algorithms. The former encompasses networks ranging from one-hop NOMA/OFDMA systems to multiple-cell and overlay systems. Meanwhile, we classify the latter into conventional QC algorithms, such as classical QAOA and quantum-inspired QPSO, and AI-driven QC algorithms like QNN, with an emphasis on comprehending their suitability and effectiveness across various network landscapes.
@@ III-B2 Applications of Conventional QC for Power Allocation
A single-cell communication system in wireless networks typically consists of a BS or access point and multiple user devices, responsible for handling wireless communications within a local area. User devices within this area communicate with the BS. A typical single-cell system is shown in Fig. 12. [116] and [117] considered power allocation within a straightforward context, focusing on a single cell of a multiple access (MA) system while applying conventional QC algorithms. $K$ transmitter-receiver communication pairs with Gaussian interference were involved in [116]. The problem was conceptualized as a non-convex and NP-hard programming formulation aimed at maximizing the weighted-sum rate. To address this problem, a QPSO based approach was employed, leveraging the characteristics of QC and biological behaviors. However, the network structure was simple and only Shannon capacity was considered without addressing interference. Energy limitation issues was addressed in [117] in cognitive radio sensor networks by constructing a mathematical model for power allocation. Bernstein approximation technique was used to control the approximation error of function value to tackle probability constraints. In addition, it adopted the concept of co-evolution to deal with large scale problems. The proposed optimization problem was complex and NP-hard, which was corresponding to the co-evolutionary theory. A stochastic optimization algorithm based on polynomial approximation and QPSO was proposed to solve the problem.
Though one-hop cooperative relay between the BS and multiple users, such as UE 1 communicated with BS via relay-UE 2 in Fig. 12, was considered [115], it was still within the area of a single-cell system. The relay link was introduced between the source node and destination node to reduce channel interference. Thus, there were two different links: the BS to relay nodes, and relay nodes to users. The strategies of power allocation for multiple relays and users significantly influenced the interference among nodes in the local communication area. The situation became worse when a full-duplex operation of channels from the perspective of MIMO was considered. Besides, another QoS metric, secrecy rates of end-users, was considered to represent the overhearing from eavesdroppers near end-users. The problem formulation was proved to be non-convex and solved by QPSO.
@@
In contrast, more complex scenarios, spanning from multiple-cell networks to overlay networks, were investigated in [118], and [119] while still utilizing conventional QC algorithms. An overlay system refers to an architecture in which additional network components or services are deployed over a single-cell system, like mesh networks, or virtualized network functions, and among others.
The problem of scheduling multiple resources including power allocation and channel assignment in a overlay OFDMA network was studied in [118]. There was an additional constraint on the channel assignment for cognitive users, which complicated the problem comparing with resource scheduling in traditional wireless cognitive networks without primary networks. The work formulated the problem of resource scheduling in a two-layer wireless network, and proposed a model to describe system outage probability according to strategies of power allocation and channel assignment. [119] discussed the resource scheduling, including power allocation and channel assignment, in a single-cell OFDMA system with $K$ users and $N$ sub-carries, in which there were two heterogeneous service requirements from users: minimum-rate guarantee services and best effort services. It can be viewed as a virtual overlay network according to two heterogeneous service requirements. The work divided the problem formulation into two optimization parts: channel assignment for a given power allocation, which was solved stochastically by QPSO based solution; power allocation for a given channel assignment, which was transformed into a concave optimization problem and solved optimally by Dinkelbach method. Finally, an acceptable solution is obtained by iterating two optimization parts. Noting that there is no analysis to obtain the optimal solution by iteration of two part even if the optimal solution is obtained.

### v3

@@ II-F Quantum-inspired Algorithms
The origin of quantum-inspired algorithms stems from the incorporation of QC concepts into classical algorithms. These algorithms are designed to run on classical computers while emulating certain quantum properties, such as superposition or entanglement, to enhance computational performance. By leveraging quantum principles, quantum-inspired algorithms offer a way to achieve improvements in efficiency and scalability, even without the use of quantum hardware.
A subclass of quantum-inspired algorithms considers their implementation on quantum hardware. For example, in QRL algorithms, state transitions are represented through measurements of an observable [80]. After the measurement, the quantum state collapses to one of the observable’s eigenstates, corresponding to the next state, and the action is probabilistically selected by the agent. While theoretically plausible, practical challenges arise in more complex scenarios. Physically implementing an observable at each step is non-trivial, and the no-cloning theorem introduces complications regarding how to generate sufficient independent copies of quantum states to perform statistical analysis on measurement outcomes.
The second subclass of quantum-inspired algorithms explicitly operates on classical hardware. A notable example is Tang’s work [81], where a binary tree-like data structure is introduced to meet quantum state preparation assumptions. This structure allows a classical algorithm to generate an $\mathcal{l}^{2}$-norm sample from a rank-$k$ approximation of a matrix in nearly polynomial time. It demonstrates how certain quantum advantages can be imitated classically under specific conditions, without requiring actual quantum computation.
The third category is referred to as quantum-inspired metaheuristic algorithms. As the name suggests, these algorithms incorporate quantum principles to modify steps in classical metaheuristic algorithms, aiming to enhance performance [82]. Numerous studies have explored this approach, including quantum ant colony algorithms [83], quantum evolutionary algorithms [84], QGA [85], and QPSO [86]. These algorithms have been applied to solve classical NP-hard problems such as the knapsack and traveling salesman problems, as well as practical issues like clustering and project scheduling.
@@
Modifications to quantum-inspired algorithms based on QC principles typically follow a consistent pattern. It involves representing original data, such as chromosomes in genetic algorithms, using qubits or n-dimensional normalized vectors. Rotation gates or matrices are then employed to evolve qubits for enhanced exploration. A random number generator is often used to generate new feasible solutions, simulating the collapse of quantum measurements to ground states. Additionally, entangled qubit states are integrated into the representation, ensuring that when a new population is created by measuring the qubit string (solution), the entangled qubits introduce diversity, which helps in avoiding local optima.
@@
Quantum-inspired metaheuristic algorithms improve optimization by leveraging QC concepts to accelerate convergence, enhance exploration and exploitation, and find optimal solutions efficiently. However, much of the existing work still lacks rigorous complexity analysis to validate these performance improvements. Another concern is the challenge of embedding QC concepts into the optimization workflow in a way that aligns with metaheuristic algorithms, without compromising the algorithm’s computational efficiency.
@@ II-G Common QC Platforms and Libraries
In the development of QC hardware, multiple technologies are being explored to realize universal quantum computers. For clarity, the key advantages and disadvantages of these hardware architectures are summarized in Table III.
TABLE III: Comparison of varied QC hardware
Hardware architecture technical route Pros Cons Challenges Typical vendors
Superconducting Use superconducting circuits to generate and manipulate qubits, employ microwave pulses to control quantum gates. Fast response, scalable, and seamlessly integrable with existing microelectronics technology. Be susceptible to environmental perturbations, including thermal noise, electromagnetic noise, etc., resulting in shorter coherent time. Maintain high fidelity while scaling. IBM, Google, Rigetti
Ion Trap Use light and magnetic fields, manipulate them with lasers. High stability, long coherence times, and high quantum gate fidelity. Cumbersome physical manipulation of ion traps, high optical stability requirements, and complex laser operating systems. Complexity in scaling up the system. IonQ
Optical QC Use photons as information carriers. Less affected by environmental noise, with relatively mature techniques for generating, manipulating, and detecting single photons. Require precise optical operations, reliable single-photon sources, and efficient single-photon detectors. Implement quantum logic gates with high efficiency and fidelity, as well as large-scale optical QC. Xanadu
Topological QC Rely on emergent quasiparticles called Majorana zero-modes, which is robust against perturbations. More stable qubits that are resistant to many types of errors and noise. Manufacture and controll Majorana zero-modes is complex and technically challenging. Reliable fabrication and detection of Majorana zero-modes in experiments have not yet been achieved. Microsoft
@@
• Qiskit [87]: Qiskit, an open-source software development kit (SDK) by IBM, provides robust tools for programming quantum computers or simulators using Python. It enables access to IBM quantum systems and offers extensive documentation with tutorials and case studies, serving as a valuable resource for both learning and research in QC. Qiskit has evolved rapidly, incorporating cutting-edge advancements such as Qiskit Machine Learning, which facilitates the creation and training of QML models. These models are capable of addressing various tasks, including classification, regression, and RL, making Qiskit a key platform for researchers and practitioners in the field.
• Cirq & Tensorflow Quantum [88]:
@@
Cirq, developed by Google’s AI quantum team, is a powerful QC framework that supports both simulation and programming for specific quantum processors. Cirq offers two distinctive components: OpenFermion, designed primarily for quantum chemistry, and TensorFlow quantum (TFQ), an open-source ML library built on TensorFlow. Cirq introduces two key data types: quantum circuit batch, which represents quantum circuits of varying sizes, and Pauli sum batch, which captures linear combinations of tensor products of Pauli operators. Using these data types, TFQ can efficiently handle tasks such as sampling the output distribution of circuit batches, computing the expected values of Pauli sums and circuit batches, calculating gradients, and simulating quantum circuits and quantum state batches. By integrating TFQ with TensorFlow, practitioners can leverage TensorFlow’s extensive capabilities, including automatic differentiation and ML tools, while also taking advantage of quantum circuit simulators and quantum hardware access. However, TFQ’s strong dependence on TensorFlow and Cirq presents compatibility challenges with other QC libraries and frameworks.
• PennyLane [89]:
@@
PennyLane is a python-based QML library that integrates classical ML algorithms with QC using automatic differentiation. It operates independently of specific underlying frameworks, allowing seamless interaction with both classical ML libraries like Pytorch and TensorFlow and QC frameworks such as Qiskit and Cirq. PennyLane uses a plugin system to connect to various quantum hardware and simulators. The library offers quantum-specific operations for evaluating and optimizing quantum circuits, encapsulates quantum circuits in quantum nodes, and supports hybrid quantum-classical ML models, including neural networks. PennyLane’s automatic differentiation quantum nodes facilitate the training of ML models and provide user-friendly interfaces for quantum algorithms and applications. However, its performance relies on the stability and quality of the plugin system, and its smaller community may result in fewer supporting resources compared to other tools integrated with traditional ML frameworks.
• Torch Quantum (QuantumNAS) [90]:
@@
Torch quantum, developed by a consortium including MIT, Yale, and the University of Texas at Austin, integrates seamlessly with Pytorch, adding several vital components: a quantum state vector simulator optimized for GPU acceleration, quantum circuit modules enabling the definition and manipulation of PQCs similar to classical neural networks, quantum optimization facilitating parameter optimization using classical algorithms like gradient descent, and an interface to real quantum devices. This extension of Pytorch allows developers to construct quantum programs using familiar tools and interfaces while benefiting from automatic differentiation for training QNNs. Additionally, the high-performance computing state vector simulator supports large-scale simulations on GPUs. However, its performance might not match specialized QC platforms like IBM quantum and Rigetti due to its reliance on Pytorch optimization.
@@ III-A Overview
Figure 12: Different network scenarios in wireless communications. From the figure, single-cell communication can be easily identified in the bottom-left, while multi-cell communication is shown on the right. Cognitive networks within multi-cell communications, featuring primary and secondary base stations (BSs)/UEs, are displayed at the top of the figure. The relay networks are displayed at the bottom.
@@ III-B1 Overview
Power allocation within wireless communication [91] involves the strategic distribution of available power resources among the diverse components of a communication system, such as BSs and user devices. The primary objective is to optimize power utilization to achieve communication that is both reliable and efficient, taking into account factors such as signal quality, network coverage, interference management, and energy consumption [92], which is also an important issue in 5G/6G cellular networks [93]. In more straightforward terms, power allocation revolves around determining the appropriate amount of power assigned to each transmitter or device in a wireless network to ensure effective communication. Power allocation is usually implemented at the physical layer, involving only nodes within one same communication cell, optimizing the transmission power of signals to maximize transmission efficiency or minimize bit error rate. For example, power optimization with relay selection have been studied [94]. By selecting appropriate relay nodes and allocating corresponding power, the energy efficiency is maximized.
TABLE IV: Summary of QC Works on Power Allocations in Wireless Networks
@@
[96] Primary users and secondary users competing for spectral resources Employ Bernstein approximation to address probability constraints and co-evolutionary theory to handle large-scale problems. Maximize network throughput QPSO Centralized
[94] One-hop cooperative relay between BS and multiple users in a NOMA system Apply QPSO to maximize QoS from the perspective of MIMO Maximize computation efficiency and energy efficiency QPSO Centralized
Overlay networks [97] Primary users and one primary BS; Cognitive users and one cognitive BS Formulate resource scheduling in a two-layer wireless network, and propose a model to describe outage probability Minimize system outage probability QPSO Centralized
[98] K users and N sub-carries in OFDMA system: two heterogeneous services, minimum-rate guarantee and best effort services Divide the problem: channel assignment given power allocation, power allocation given channel assignment; transform it into a concave optimization and solved by Dinkelbach method Maximize energy efficiency QPSO Centralized
@@
TABLE V: Details of QC Algorithms for Power Allocation
@@
In this section, we undertake an investigation of existing QC-driven approaches along two fundamental dimensions: the complexity of network scenarios and the categories of QC algorithms. The former encompasses networks ranging from one-hop non-orthogonal channel multiple access (NOMA)/orthogonal frequency-division multiple access (OFDMA) systems to multiple-cell and overlay systems. Meanwhile, we classify the latter into conventional QC algorithms, such as classical QAOA and quantum-inspired QPSO, and AI-driven QC algorithms like QNN, with an emphasis on comprehending their suitability and effectiveness across various network landscapes.
@@ III-B2 Applications of Conventional QC for Power Allocation
A single-cell communication system in wireless networks typically consists of a BS or access point and multiple user devices, responsible for handling wireless communications within a local area. User devices within this area communicate with the BS. A typical single-cell system is shown in Fig. 12. [95] and [96] consider power allocation within a straightforward context, focusing on a single cell of a multiple access (MA) system while applying conventional QC algorithms. $K$ transmitter-receiver communication pairs with Gaussian interference are involved in [95]. The problem is conceptualized as a non-convex and NP-hard programming formulation aimed at maximizing the weighted-sum rate. To address this problem, a QPSO based approach is employed, leveraging the characteristics of QC and biological behaviors. However, the network structure is simple and only Shannon capacity is considered without addressing interference. Energy limitation issues are addressed in [96] in cognitive radio sensor networks by constructing a mathematical model for power allocation. Bernstein approximation technique is used to control the approximation error of function value to tackle probability constraints. In addition, it adopts the concept of co-evolution to deal with large scale problems. The proposed optimization problem is complex and NP-hard, which is corresponding to the co-evolutionary theory. A stochastic optimization algorithm based on polynomial approximation and QPSO is proposed to solve the problem.
Though one-hop cooperative relay between the BS and multiple users, such as UE 1 communicated with BS via relay-UE 2 in Fig. 12, is considered [94], it is still within the area of a single-cell system. The relay link is introduced between the source node and destination node to reduce channel interference. Thus, there are two different links: the BS to relay nodes, and relay nodes to users. The strategies of power allocation for multiple relays and users significantly influence the interference among nodes in the local communication area. The situation becomes worse when a full-duplex operation of channels from the perspective of multiple input multiple output antennas (MIMO) is considered. Besides, another QoS metric, secrecy rates of end-users, is considered to represent the overhearing from eavesdroppers near end-users. The problem formulation is proved to be non-convex and solved by QPSO.
@@
In contrast, more complex scenarios, spanning from multiple-cell networks to overlay networks, are investigated in [97], and [98] while still utilizing conventional QC algorithms. An overlay system refers to an architecture in which additional network components or services are deployed over a single-cell system, like mesh networks, or virtualized network functions, and among others.
The problem of scheduling multiple resources including power allocation and channel assignment in an overlay OFDMA network is studied in [97]. There is an additional constraint on the channel assignment for cognitive users, which complicates the problem comparing with resource scheduling in traditional wireless cognitive networks without primary networks. The work formulates the problem of resource scheduling in a two-layer wireless network, and proposes a model to describe system outage probability according to strategies of power allocation and channel assignment. [98] discusses the resource scheduling, including power allocation and channel assignment, in a single-cell OFDMA system with $K$ users and $N$ sub-carries, in which there are two heterogeneous service requirements from users: minimum-rate guarantee services and best effort services. It can be viewed as a virtual overlay network according to two heterogeneous service requirements. The work divides the problem formulation into two optimization parts: channel assignment for a given power allocation, which is solved stochastically by QPSO based solution; power allocation for a given channel assignment, which is transformed into a concave optimization problem and solved optimally by Dinkelbach method. Finally, an acceptable solution is obtained by iterating two optimization parts. Noting that there is no analysis to obtain the optimal solution by iterations even if the optimal solutions for two parts are obtained. Specifically from the QC perspective, there are three modules same with the traditional PSO algorithm: particle position, fitness function, and evolution equation. The difference lies in that the iterative equation in the evolution operation for the mean best position bits, namely, decision variables of assigning frequency resource units to multiple users, is determined by the Hamming distance of two input binary bits. The $j$-th bit in the mean best solution from all particles is determined by the states of the corresponding $j$-th bit of each best particle.